find the remainder when polynomial 3x^100+5x^85-4x^38+2x^17-6 is divided by x+1
Answer by stanbon(75887)

find the remainder when polynomial 3x^100+5x^85-4x^38+2x^17-6 is divided by x+1
------
Find f(-1) = 3*1+5*-1-4*1+2*-1-6 =3-5-4-2-6 = -2
The remainder is -2
======================================================
Cheers,
Stan H.